1)Amoung 18 members of a cricket club, there are 2 wicket keepers, 5 bowlers and 6 batsman and 5 allrounders. What is the probability that a randomly selected team of 12 players comprise of 1 wicketkeeper, 3 bowlers, 4 batsman and remaining allrounders
Answer by stanbon(75887)

1)Among 18 members of a cricket club, there are 2 wicket keepers, 5 bowlers and 6 batsman and 5 allrounders. What is the probability that a randomly selected team of 12 players comprise of 1 wicketkeeper, 3 bowlers, 4 batsman and remaining allrounders
------------------
If there are 12 members on a cricket team,
Ways to succeed: 2C1 * 5C3 *6C4 * 5C4 = 2*10*15*5 = 1500
------
Ways to select 12 from 18 = 18564
======================================
Probability = 1500/18564 = 0.0808
======================================
Cheers,
Stan H.